# Zero-shot Class Unlearning via Layer-wise Relevance Analysis and Neuronal Path Perturbation

### v1

@@ Abstract
In the rapid advancement of artificial intelligence, privacy protection has become crucial, giving rise to machine unlearning. Machine unlearning is a technique that removes specific data influences from trained models without the need for extensive retraining. However, it faces several key challenges, including accurately implementing unlearning, ensuring privacy protection during the unlearning process, and achieving effective unlearning without significantly compromising model performance. This paper presents a novel approach to machine unlearning by employing Layer-wise Relevance Analysis and Neuronal Path Perturbation. We address three primary challenges: the lack of detailed unlearning principles, privacy guarantees in zero-shot unlearning scenario, and the balance between unlearning effectiveness and model utility. Our method balances machine unlearning performance and model utility by identifying and perturbing highly relevant neurons, thereby achieving effective unlearning. By using data not present in the original training set during the unlearning process, we satisfy the zero-shot unlearning scenario and ensure robust privacy protection. Experimental results demonstrate that our approach effectively removes targeted data from the target unlearning model while maintaining the model’s utility, offering a practical solution for privacy-preserving machine learning.
@@
Machine unlearning, privacy preservation, model security.
@@ I Introduction
In recent years, with the growing public attention on AI models, research has increasingly focused on privacy concerns stemming from user-model interactions. Once a model is trained, users may wish to withdraw their data and remove its influence from the model. Traditionally, this can be achieved by retraining the model, but retraining is time-consuming and inefficient for large datasets, such as those used in large language models (LLMs) [1]. The goal of unlearning is to address this issue by selectively removing the model’s memory of specific data without retraining the entire model. This can be done through data reorganization or model manipulation, allowing compliance with data deletion requests and safeguarding user privacy [2].
Machine unlearning research is primarily divided into class unlearning and sample unlearning [2, 3]. Class unlearning aims to eliminate all knowledge the model has acquired from a specific data class. This type of unlearning is often necessary when data usage policies require the deletion of certain users’ data, when the data has lost its relevance, or when it has introduced bias. In contrast, sample unlearning focuses on removing the influence of specific data samples from the model. This approach is commonly applied when users request the deletion of their data or when problematic data samples need to be excluded. Class unlearning is better suited for deleting data related to broad groups, while sample unlearning is better for individual user needs. By exploring practical ways to implement class and sample unlearning, researchers can develop more flexible privacy protection mechanisms that meet user demands without incurring unnecessary performance losses.
So far, class unlearning is particularly important and requires more research because it tackles the challenge of removing knowledge from entire categories of data, which is crucial for mitigating risks such as data misuse and unintended bias. Class unlearning methods have been applied to a wide range of AI tasks. For example, Chen et al. [4] achieved unlearning in image classification models by manipulating the model’s decision space. Ma et al. [5] implemented unlearning using neuron masks derived from gradients. Furthermore, Sun et al. [6] employed unlearning methods to make StyleGAN forget specific knowledge classes. Additionally, Chen et al. [7] proposed an unlearning framework for Graph Neural Networks (GNNs). These examples illustrate that class unlearning goes beyond simply removing individual samples—it ensures that entire problematic or sensitive data categories can be eliminated, which is essential for building fair and privacy-preserving AI systems.
Class unlearning can be approached using data reorganization and model manipulation [2]. However, both methods face a critical problem in balancing effective unlearning with maintaining model utility. The instability of the unlearning process presents challenges in two primary forms: over-unlearning and under-unlearning. Over-unlearning occurs when excessive data is removed, diminishing the model’s utility. At the same time, under-unlearning happens when insufficient data is erased, preserving the model’s usefulness but failing to achieve the unlearning objective.
@@
  • Lack of explainable machine unlearning methods. A prominent challenge in machine unlearning is the need for more basic explanations regarding the unlearning principle. Even though users submit unlearning requests to model owners, there’s still a significant lack of clear explanations on how the unlearning methods influence the models’ parameters. This lack of clarity about unlearning principles makes it hard for users to understand how unlearning works, preventing them from trusting these methods.
@@
  • Model owners must access users’ original data to use unlearning methods. Another critical challenge confronting machine unlearning is more privacy guarantees. Many existing methods use the users’ original model for unlearning, either by fine-tuning or by directly modifying the parameters. Using original data raises serious privacy concerns because it may expose sensitive information during model updates. Focusing on data-driven methods worsens the privacy problem, highlighting the urgent need for zero-shot machine unlearning methods.
  • The time and computational resources needed to perform machine unlearning are still excessively high. Many existing unlearning methods necessitate extensive calculations, which can involve retraining or fine-tuning models based on large datasets. This high resource demand increases operational costs and introduces delays, making it less feasible for real-world applications where efficiency and responsiveness are critical. As a result, achieving practical implementation remains a considerable hurdle.
@@
We address the three challenges mentioned above using different approaches, thereby designing an efficient unlearning method. We investigate the relevance between the model’s neurons and the unlearning information for the first challenge. ”Relevance” refers to the contribution or importance of each neuron in the neural network to the final prediction. We identify the path of neurons with high relevance, which we define as the classification path. Our goal is to determine the neuron classification path by assessing the relevance of each neuron to the unlearning information. Through layer-wise relevance analysis, we can clearly describe the internal principles of machine unlearning.
For the second challenge, we use data not present in the training dataset to perform actions in the unlearning process, such as queries. This kind of data can be obtained in two ways. First, we can directly extract the data from the test set. When the model training is complete, the model owner will conduct performance evaluations using data that was not used for training, which can also be used for machine unlearning. Additionally, data generated by generative models can be utilized in the unlearning process. We can instruct generative models to generate data related to unlearning information and use queries or other methods to determine the relevance of neurons in the model. Through these two methods, we can ensure the privacy of users’ data during the machine unlearning process.
One last point is that our method does not require modifying the overall model parameters after the model processes the data, significantly saving time and computational resources. We analyze the data related to the unlearning information and identify the neurons that show high relevance. Due to the redundancy in neural networks, we can achieve effective unlearning by perturbing only a small subset of neurons, which minimizes the need for time and computational resources. This approach enhances the feasibility of real-world applications, enabling efficient unlearning while maintaining model utility.
@@
  • We propose an explainable unlearning method significantly improving unlearning speed and reducing resource consumption. By employing layer-wise relevance analysis and neuronal path perturbation, we can adjust the neuron weights highly related to unlearning information within the model. Thereby, we can achieve effective unlearning without compromising the model’s utility.
@@
  • We demonstrate that unlearning in neural networks can be approached from the perspective of network inference by conducting experiments and analyses focused on model parameters and redundancy. Our method addresses traditional parameter adjustments and delves into the redundant characteristics within neural networks. This inference-based perspective offers new insights and methodologies for future research.
@@ II-A Machine Unlearning
First, Bourtoule et al. [8] developed an unlearning algorithm where users find it difficult to revoke access to their data and request its deletion after sharing. They introduced the SISA training framework, which speeds up the unlearning process by strategically limiting the impact of data points during training. Chen et al. [4] proposed Boundary Unlearning, which rapidly unlearns an entire class from a DNN model by shifting its decision boundary rather than altering the model parameters. Jia et al. [9] introduced an approach to machine unlearning by shifting focus from data-centric methods to a model-based perspective, specifically using model sparsification through weight pruning. This ”prune first, then unlearn” paradigm reduces the gap between exact and approximate unlearning. Chen et al. [10] proposed a fast model debiasing framework (FMD) that identifies and removes biases in trained models using counterfactual concepts and influence functions. They integrate a machine unlearning-based strategy to efficiently debias models with minimal data and parameter updates. Kurmanji et al. [11] introduced SCRUB, a novel unlearning algorithm that excels across different unlearning applications, including removing biases (RB), resolving confusion (RC) and protecting user privacy (UP). Foster et al. [12] proposed Selective Synaptic Dampening (SSD), a novel retrain-free machine unlearning method that uses the Fisher information matrix to identify and dampen parameters critical to the data being forgotten. SSD does not require long-term training data storage and achieves performance comparable to retrain-based methods. Liu et al. [13] explored the security risks of machine unlearning by introducing backdoor attacks exploiting unlearning requests. They propose two attack methods: one where the attacker submits unlearning requests without poisoning any training data and another where the attacker poisons a few data instances with a pre-defined trigger and then activates the attack through malicious unlearning.
@@
Moreover, Chundawat et al [14] addressed the challenge of achieving machine unlearning in scenarios without original training data, introducing the concept of zero-shot machine unlearning. They propose two methods, error minimizing-maximizing noise and gated knowledge transfer, to remove the influence of the forgotten data while preserving model performance on retained data. Tarun et al. [15] introduced a fast and scalable machine unlearning framework that uses error-maximizing noise and weight manipulation to unlearn specific data classes without accessing the entire training dataset. Their method applies an impair-repair approach, where noise is used to sharply unlearn a target class, followed by a repair step to recover overall model performance. Existing research methods have also extended to feature unlearning. Xu et al. [16] introduced a ”feature unlearning” scheme to address the limitations of existing machine unlearning methods, which typically focus on removing entire instances or classes. Recognizing the need for feature-level unlearning to maintain model utility They further proposed an approach [17] constructed an essential graph to describe relationships between critical parameters, identify parameters important to both the forgotten and retained targets and use a pruning-based technique to remove information about the target to be unlearned effectively.
Compared to the abovementioned methods, our approach focuses more on the correlation between neurons in the neural network and the unlearning data. Additionally, our method considers user privacy, as it does not require using the user’s original model, which has not received much attention in previous methods. Ultimately, our approach efficiently accomplishes the unlearning task while maintaining the model’s utility.
@@ II-B Layer-wise Relevance Propagation
Today, research on LRP still holds a significant influence. Li et al. [20] proposed the Weight-dependent Baseline Layer-wise Relevance Propagation algorithm, which improves instance-level explanations by addressing limitations in existing LRP methods that disregard model weights or sample features. Bassi et al. [21] proposed ISNet, which optimizes LRP heatmaps to reduce background bias and mitigate shortcut learning in deep classifiers without increasing computational cost. Achtibat et al. [22] extended Layer-wise Relevance Propagation to attention layers in transformer models, providing a computationally efficient method for faithfully attributing input and latent representations. Hatefi et al. [23] proposed optimizing the hyperparameters of attribution methods from eXplainable AI for network pruning, demonstrating improved model compression rates for both transformer-based and convolutional networks.
@@ III-B Layer-wise Relevance Propagation
Layer-wise Relevance Propagation (LRP) is a technique used to interpret the decisions made by deep neural networks. Let $F(x)$ represent the model’s prediction function for an input $x$. Given a neural network, LRP aims to attribute the prediction $F(x)$ to the individual input features by propagating relevance scores through the network layers.
@@
The relevance scores are propagated from the output layer back to the input layer using a set of propagation rules. Let $R_{k}^{(l_{i})}$ be the relevance score of neuron $k$ in layer $l_{i}$, and we denote all the different neurons in layer $l_{i}$ as $k^{\prime}$. The relevance score is propagated to the layer $l_{i}$ from layer $l_{i+1}$ according to the following general rule:
$R_{k}^{i}=\sum_{j}\frac{z_{kj}}{\sum_{k^{\prime}}z_{k^{\prime}j}}R_{j}^{i+1},\ k\in l_{i},\ k^{\prime}\in l_{i},\ j\in l_{i+1}$ (4)
@@ IV-A Threat Model
Users: Users can know that their data is used to train the model and might be aware of privacy policies and data usage agreements. They can request the unlearning of their data and withdraw consent for its use in the model. After requesting machine unlearning, users may receive confirmation that their data has been removed.
Model owner: Model owners know the details of the training dataset, including data provided by each user, and understand the model’s performance metrics and how unlearning affects it. They can locate and delete specific user data upon request, retrain or adjust the model to maintain its performance and ensure compliance with relevant laws. They may need to provide users with proof of unlearning.
@@ IV-B Overview of Our Method
Our method includes several steps to achieve machine unlearning. In Figure 1, after users send unlearning requests, the model owner will prepare dataset $D^{\prime}$ for the following actions. The two primary steps are layer-wise relevance analysis and neuronal path perturbation.
Layer-wise Relevance Analysis obtain and analyze all the neurons’ relevance to the classification via LRP. According to LRP rules $r$, the predictions’ attributions to the model’s inputs are obtained through backward propagation among neurons. Therefore, we can determine the relevance of neurons to classification by comparing the relevance scores $R$ of neurons in each layer. As shown in Figure 1, it is necessary to identify the neurons’ occurrence frequency after analyzing each sample due to the unique characteristics of the samples.
Neuronal Path Perturbation perturbing the neurons in the intersection of relevant neurons to achieve unlearning. We can conduct perturbation just for the neurons in the fully connected layer before the output layer. If only the output layer is perturbed, it is similar to abandoning the model’s output of a specific class. Similarly, in Figure 1, we first confirm the target neurons quantity $m_{p}$. Then, we can perturb the target neurons to cut off the model’s neuronal classification path to achieve machine unlearning.
After these two steps, we can get the class-unlearned model without additional training. We will describe our method in the following section.
@@ IV-C Layer-wise Relevance Analysis
A neuron network can be seen as a complex function that can finish a specific task. The networks’ predictions are the results of neurons’ internal interactions. We need to understand the internal principle of neuron networks to achieve unlearning without additional training. However, obtaining the relevance of neurons to the model’s inputs is difficult due to neuron networks’ black-box characteristics.
@@
To fix this problem, we first apply layer-wise relevance propagation (LRP) to obtain each neuron’s internal relevance score $R$. We prepare a dataset $D^{\prime}$ that contains no users’ data to conduct LRP. The relevance $R$ is conserved through the layers, ensuring that the total relevance is preserved from the output layer back to the input layer. This process can be demonstrate in Equation 5 and 6:
$R_{j}^{l+1}=\sum_{i}R_{i\leftarrow j}^{l,l+1}$ (5)
@@
Here, based on LRP, $R_{j}^{l+1}$ is the relevance at layer $l+1$ for neuron $j$, and $R_{i\leftarrow j}^{l,l+1}$ is the relevance propagated from neuron $j$ at layer $l+1$ to neuron $i$ at layer $l$. For layer $l$, the relevance of neuron $i$ can be computed by summing all the relevance from neurons at layer $l+1$.
$R_{i}^{l}=\sum_{j\in(l+1)}R_{i\leftarrow j}^{l,l+1}$ (6)
Via the method above, we can calculate the relevance of each neuron about input $x$ in the neural network. For the FC layer $l_{i}$, we pair the neuron positions in this layer with their relevance and sort them according to the order of the neuron positions. Assuming there are $m$ neurons in layer $l_{i}$ and the position of neuron $j$ is $p_{j}^{i}$ we can demonstrate this step as:
$List(l_{i},x)=Sort\{(R_{0}^{i},p_{0}^{i}),...,(R_{j}^{i},p_{j}^{i}),...,(R_{m-1}^{i},p_{m-1}^{i})\},$ (7)
$p_{0}^{i}<p_{1}^{i}<\ldots<p_{m-1}^{i}$
@@
We define the top $k$ neurons with high $R$s in $List(l_{i},x)$ as high-relevance neurons. By extracting these top $k$ neurons in $List(l_{i},x)$, we obtain $List(l_{i},x,k)$ as Equation 8. Unlike Equation 7, the $(R,p)$ pairs in $List(l_{i},x,k)$ are arranged in descending order of $R$, rather than by the position of the neurons in $l_{i}$.
$List(l_{i},x,k)=\{(R_{0}^{i},p_{0}^{i}),(R_{1}^{i},p_{1}^{i}),\ldots,(R_{k-1}^{i},p_{k-1}^{i})\},$ (8)
$R_{0}^{i}\geq R_{1}^{i}\geq\ldots\geq R_{k-1}^{i},k\in\left[1,m\right]$
@@
However, the action above only conducts relevance analysis for a single sample. We need to analyze more data to identify a set of neurons generally relevant to the data within an unlearning dataset. We use dataset $D^{\prime}$ for our analysis, which includes data from the unlearning class. These data are extracted from open-source datasets or generated by generative models to protect user privacy. We define that dataset $D^{\prime}$ contains $n$ data samples. After conducting relevance analysis in layer $l_{i}$ on each sample, we calculate the occurrence frequency $C$ of neurons to eliminate the influence of data-specific characteristics.
$C\left(p^{i}\right)=\sum_{j\in n}{Count}\left(List\left(l_{i},x_{j},k\right)\right),x_{j}\in D^{\prime},k\in\left[1,m\right]$ (9)
As shown in Equation 9, we define $C(p)$ as the occurrence frequency of neuron $p^{i}$ in all the $List(l_{i},x,k)$, which is calculated using the counting function $Count$.
@@
For all neurons with high $R$, we obtain their occurrence frequency $C$ using the method mentioned above. Next, as in Equation 10, we sort these $C$ values in descending order to identify the strongly relevant neurons $S^{i}$ corresponding to unlearned class in layer $l_{i}$.
@@
It is worth mentioning that due to the specificity of the images, the neurons associated with each data sample will vary. Therefore, if we denote the number of neurons that appear in all the lists as $z$, the value of z will range between $k$ and $m$.
$S^{i}=Sort\left\{C\left(p_{0}^{i}\right),C\left(p_{1}^{i}\right),...C\left(p_{z-1}^{i}\right)\right\},$ (10)
$p_{0}^{i}<p_{1}^{i}<\ldots<p_{z-1}^{i},z\in\left[k,m\right]$

### v2

@@ Abstract
Machine unlearning is a technique that removes specific data influences from trained models without the need for extensive retraining. However, it faces several key challenges, including the lack of explanation, privacy concerns during the unlearning process, and the high demand for time and computational resources. This paper presents a novel unlearning approach to tackle above challenges by employing Layer-wise Relevance Analysis and Neuronal Path Perturbation. Our method balances machine unlearning performance and model utility by identifying and perturbing highly relevant neurons, thus achieving effective unlearning. Using unseen data that has not been presented in the original training set, our method achieves zero-shot unlearning, which allows for the removal of specific class knowledge without accessing the original training data during the unlearning process. This approach ensures robust privacy protection. Experimental results demonstrate that our approach effectively removes targeted data from the target unlearning model while maintaining the model’s utility, offering a practical solution for privacy-preserving machine learning.
@@
Machine unlearning, privacy preservation, neuronal path perturbation.
@@ I Introduction
In recent years, with the growing public attention on Artificial Intelligence (AI) models, research has increasingly focused on privacy concerns stemming from user-model interactions. Once a model is trained, users may wish to withdraw their data and remove its influence from the model. Traditionally, this can be achieved by retraining the model, but retraining is time consuming and inefficient for large datasets, such as those used in large language models (LLM) [1]. The goal of unlearning is to address this issue by selectively removing the model’s memory of specific data without retraining the entire model. This can be done through data reorganization or model manipulation, allowing compliance with data deletion requests and safeguarding user privacy [2, 3].
Machine unlearning research is primarily divided into class unlearning and sample unlearning [2, 4]. Class unlearning aims to eliminate all knowledge the model has acquired from a specific data class. This type of unlearning is often necessary when data usage policies require the deletion of certain users’ data, when the data has lost its relevance, or when it has introduced bias. In contrast, sample unlearning focuses on removing the influence of specific data samples from the model. This approach is commonly applied when users request the deletion of their data or when problematic data samples need to be excluded.
Class unlearning methods have been applied to a wide range of AI tasks. For example, Chen et al. [5] achieved unlearning in image classification models by manipulating the model’s decision space. Ma et al. [6] implemented unlearning using neuron masks derived from gradients. Furthermore, Sun et al. [7] employed unlearning methods to make StyleGAN forget specific knowledge classes. Additionally, Chen et al. [8] proposed an unlearning framework for Graph Neural Networks (GNNs). These examples illustrate that class unlearning goes beyond simply removing individual samples—it ensures that entire problematic or sensitive data categories can be eliminated, which is essential for building fair and privacy-preserving AI systems.
@@
  • Lack of explainable machine unlearning methods. A prominent challenge in machine unlearning is the need for more basic explanations regarding the unlearning principle. Even though users submit unlearning requests to model owners, there’s still a significant lack of clear explanations on how the unlearning methods influence the models’ parameters.
@@
  • Model owners may need to access users’ original data to use unlearning methods. Many existing methods use the users’ original data for unlearning, either by fine-tuning or by directly modifying the parameters. Using original data raises serious privacy concerns because it may expose sensitive information during model updates.
  • High cost on time and computational resources Many existing unlearning methods necessitate extensive calculations, which can involve retraining or fine-tuning models based on large datasets.
@@
We address the three challenges by investigating the relevance between the model’s neurons and the unlearning information.“Relevance” refers to the contribution or importance of each neuron in the neural network to the final prediction. We identify the path of neurons with high relevance, which we define as the classification path. Our goal is to determine the neuron classification path by assessing the relevance of each neuron to the unlearning information. Through layer-wise relevance analysis, we can explain the internal principles of machine unlearning.
For the second challenge, our solution follows the concept of zero-shot machine unlearning as defined by Chundawat et al. [9]. In their work, zero-shot unlearning refers to performing the unlearning process without needing access to the original training dataset (including retain data and forget data), rather than not using any data. Based on this definition, our method utilizes “unseen samples”, which are data not present in the original training dataset, to perform the unlearning process. These unseen samples can be either directly extracted from the test set or generated by generative models. By relying solely on such data, we effectively adhere to the zero-shot unlearning paradigm and help ensure the privacy of users’ original training data during the machine unlearning process.
In terms of the method’s implementation efficiency, the proposed method does not need to modify the overall model parameters, significantly mitigate the time and computational cost. We analyze the data related to the unlearning class information and identify the neurons that show high relevance. Based on this, our method applies Neuronal Path Perturbation by selectively disconnecting these neurons within a single fully connected layer from their connections to preceding and succeeding layers. This action perturbs the model’s classification path, causing it to lose the ability to classify the unlearning class data. Due to the inherent redundancy in neural network paths for other classes and our perturbation being limited to a single layer, the model’s overall utility is largely preserved.
@@
  • We propose an explainable zero-shot unlearning method significantly improving unlearning speed and reducing resource consumption. Using layer-wise relevance analysis and neuronal path perturbation, we can adjust neuron weights highly related to unlearning information within the model. Thereby, we can achieve effective unlearning without compromising the model’s utility.
@@
  • We demonstrate that unlearning in neural networks can be approached from the perspective of network inference by conducting experiments. Our method addresses traditional parameter adjustments and delves into the redundant characteristics within neural networks.
@@ II-A Machine Unlearning
First, Bourtoule et al. [10] introduced the SISA training framework, which accelerates unlearning by limiting the impact of data points during training, making it easier for users to revoke data access and request deletion after sharing. Chen et al. [5] developed Boundary Unlearning, which removes a class from a DNN by shifting its decision boundary, enabling rapid unlearning. Jia et al. [11] proposed a model-based approach to unlearning, using weight pruning for model sparsification in a ”prune first, then unlearn” method. Researchers can also use unlearning methods to remove model bias. Chen et al. [12] introduced the FMD framework, which identifies and removes biases from trained models using counterfactual concepts and influence functions. Kurmanji et al. [13] proposed SCRUB, a versatile unlearning algorithm for removing biases, resolving confusion, and protecting user privacy across various applications. Foster et al. [14] introduced Selective Synaptic Dampening (SSD), a retrain-free unlearning method that uses the Fisher information matrix to dampen parameters critical to forgotten data. Liu et al. [15] explored security risks in unlearning, proposing two backdoor attack methods: one exploiting unlearning requests without data poisoning, and another involving poisoned data with a trigger activated by malicious unlearning.
@@
However, most unlearning methods require the use of original training data, which increases the frequency of privacy data exposure and the risk of leakage. As a result, researchers have begun to focus on unlearning methods that do not rely on the training data. For example, Chundawat et al [9] introduced zero-shot machine unlearning, tackling unlearning without access to original training data. They propose two methods—error-minimizing noise and gated knowledge transfer—to remove the influence of forgotten data while preserving model performance on retained data. Tarun et al. [16] presented a fast, scalable unlearning framework using error-maximizing noise and weight manipulation. Their impair-repair approach unlearns specific data classes without the full training dataset, followed by a repair step to recover model performance. Existing research methods have also extended to feature unlearning.
Beyond the class or sample level, machine unlearning research is expanding. Researchers are now also conducting more thorough investigations at the feature level, seeking more detailed privacy protection. Xu et al. [17] introduced a “feature unlearning” scheme to address the limitations of existing machine unlearning methods, which typically focus on removing entire instances or classes. Recognizing the need for feature-level unlearning to maintain model utility. They further proposed an approach [18] constructed an essential graph to describe relationships between critical parameters, identify parameters important to both the forgotten and retained targets and use a pruning-based technique to remove information about the target to be unlearned effectively.
Our approach differs significantly by operating at the neuron level, focusing on the relevance of individual neurons to the target class to be unlearned. This enables a highly fine-grained and precise unlearning process, thereby enhancing its overall granularity. Furthermore, our method prioritizes user privacy by not requiring access to the original training data, effectively achieving zero-shot unlearning. This distinct approach ensures efficient unlearning while effectively maintaining model utility.
@@ II-B Layer-wise Relevance Propagation
Today, research on LRP still holds a significant influence. Li et al. [21] proposed the Weight-dependent Baseline Layer-wise Relevance Propagation algorithm, which improves instance-level explanations by addressing limitations in existing LRP methods that disregard model weights or sample features. Hatefi et al. [22] proposed optimizing the hyperparameters of attribution methods from eXplainable AI for network pruning, demonstrating improved model compression rates for both transformer-based and convolutional networks.
@@ III-B Layer-wise Relevance Propagation
Layer-wise Relevance Propagation is a technique used to interpret the decisions made by deep neural networks. Let $F(x)$ represent the model’s prediction function for an input $x$. Given a neural network, LRP aims to attribute the prediction $F(x)$ to the individual input features by propagating relevance scores through the network layers.
@@
The relevance scores are propagated from the output layer back to the input layer using a set of propagation rules. Let $R_{k}^{l}$ be the relevance score of neuron $k$ in layer $l$, and we denote all the different neurons in layer $l$ as $k^{\prime}$. The relevance score is propagated from layer $l+1$ to the layer $l$ according to the following general rule:
$R_{k}^{l}=\sum_{j}\frac{z_{kj}}{\sum_{k^{\prime}}z_{k^{\prime}j}}R_{j}^{l+1},\ k\in l,\ k^{\prime}\in l,\ j\in l+1$ (4)
@@ IV-A Threat Model
Users: Users can know that their data is used to train the model and might be aware of privacy policies and data usage agreements. They can request the unlearning of that specific class and withdraw consent for its use in the model. After requesting class unlearning, users may receive confirmation that the model’s knowledge of data from that class has been removed.
Model owner: Model owners cannot directly access user data and only have insight into the class distribution of the training dataset. They understand the model’s performance metrics and how unlearning affects it. They can locate and delete specific user data upon request, retrain, or adjust the model to maintain its performance, and ensure compliance with relevant laws. They may need to provide users with proof of unlearning.
@@ IV-B Overview of Our Method
Our method for achieving machine unlearning comprises several steps. As illustrated in Figure 1, after receiving unlearning requests from users, the model owner prepares dataset $D^{\prime}$, extracted from the test set, for subsequent processing. The two primary steps involved are Layer-wise Relevance Analysis and Neuronal Path Perturbation.
Layer-wise Relevance Analysis. This step obtains and analyzes the relevance of all neurons to the classification task using Layer-wise Relevance Propagation. According to LRP rules $r$, the contributions of the model’s inputs to the predictions are determined through backward propagation through the network layers. By comparing the relevance scores $R$ of neurons in each layer, we can assess their importance to the classification. As depicted in Figure 1, analyzing the relevance for each unlearning sample allows us to identify neurons that are frequently highly relevant or critical for the target classification, especially considering the specific characteristics of the samples to be unlearned.
Neuronal Path Perturbation. This step involves perturbing selected neurons, particularly those identified as relevant or critical, to achieve unlearning. Perturbation can be applied, for example, to neurons in the fully connected layer preceding the output layer. Perturbing only the output layer would be akin to simply suppressing the model’s output for a specific class, rather than truly unlearning. As illustrated in Figure 1, we first determine the quantity $m_{p}$ of target neurons to be perturbed. We then perturb these neurons to disrupt or sever the model’s neuronal pathways responsible for the target classification, thereby achieving machine unlearning.
Following these two steps, a class-unlearned model is obtained without requiring additional training. The details of our method are described in the following section.
@@ IV-C Layer-wise Relevance Analysis
A neural network can be viewed as a complex function that performs a specific task. Network predictions result from internal interactions among its neurons. To achieve unlearning without additional training, we need to understand the internal workings of neural networks. However, obtaining the relevance of neurons to the model’s inputs is challenging due to the black-box nature of neural networks.
@@
To address this challenge, we first apply LRP to obtain an internal relevance score $R$ for each neuron. We utilize dataset $D^{\prime}$ to conduct the relevance analysis. While LRP is typically applied to inputs, here we use $D^{\prime}$ (which contains data related to the unlearning class) to analyze relevance patterns across multiple relevant samples. The relevance $R$ is conserved through the layers during backpropagation, ensuring that the total relevance is preserved from the output layer back to the input layer. This backward propagation process can be described by Equation 5:
[…]
@@
Here, based on LRP, $R_{i}^{l}$ is the relevance assigned to neuron $i$ at layer $l$, and $R_{i\leftarrow j}^{l,l+1}$ is the portion of relevance propagated from neuron $j$ at layer $l+1$ to neuron $i$ at layer $l$. For layer $l$, the total relevance of neuron $i$ ($R_{i}^{l}$) is computed by summing all relevance propagated to it from neurons at layer $l+1$.
Using this method, we calculate the relevance $R_{j}^{l}$ for each neuron $j$ in layer $l$ with respect to a given input $x$. For layer $l$, assuming there are $m$ neurons with positions $p_{0},\dots,p_{m-1}$ such that $p_{0}<p_{1}<\dots<p_{m-1}$, we create a list of pairs $(R_{j}^{l},p_{j})$ ordered by neuron position. We represent this list as:
[…]
@@
We identify the top $k$ neurons with the highest relevance scores $R$ in $List(l,x)$ as high-relevance neurons for input $x$. By extracting these top $k$ pairs and sorting them in descending order of $R$, we obtain $List(l,x,k)$ as defined in Equation 7. Unlike Equation 6, the pairs $(R,p)$ in $List(l,x,k)$ are arranged in descending order of $R$, rather than by neuron position.
$List(l,x,k)=((R_{0}^{l},p_{0}),\ldots,(R_{k-1}^{l},p_{k-1})),$ (7)
$\text{where }R_{0}^{l}\geq R_{1}^{l}\geq\ldots\geq R_{k-1}^{l},k\in\left[1,m\right]$
@@
However, the analysis described above is performed for a single input sample $x$. To identify a set of neurons consistently relevant across the data associated with the unlearning class, we analyze dataset $D^{\prime}$. As mentioned, $D^{\prime}$ includes data samples from the unlearning class, either extracted from open-source datasets or generated synthetically to protect user privacy. Let $D^{\prime}=\{x_{1},\ldots,x_{n}\}$ be the set of $n$ data samples in $D^{\prime}$. After conducting the top-$k$ relevance analysis for layer $l$ on each sample $x_{j}\in D^{\prime}$, we calculate the occurrence frequency of each neuron position to mitigate the influence of sample-specific variations.
$C(p)=\sum_{j=1}^{n}\mathbb{I}\left(p\in\{p\mid(R,p)\in List(l,x_{j},k)\}\right)$ (8)
As defined in Equation 8, $C(p)$ is the occurrence frequency of a specific neuron position $p$ across all $n$ top-$k$ relevance lists $List(l,x_{j},k)$ for samples $x_{j}\in D^{\prime}$. The indicator function $\mathbb{I}(\cdot)$ evaluates to 1 if the condition is true (i.e., if neuron position $p$ is present in the list of top-$k$ neuron positions for sample $x_{j}$) and 0 otherwise.
@@
For all unique neuron positions that appeared in any of the top-$k$ lists (i.e., positions $p$ with $C(p)>0$), we obtain their occurrence frequency $C(p)$ using the method described above. Let $P_{unique}=\{p\mid C(p)>0\}$ be the set of these $z=|P_{unique}|$ unique neuron positions. Next, as represented in Equation 9, we sort these neuron positions based on their frequencies $C(p)$ in descending order to identify the list of strongly relevant neurons $S_{l}$ for the unlearned class in layer $l$.
@@
It is worth mentioning that due to sample-specific variations in relevance, the set of top-$k$ relevant neurons can vary across different samples. If we denote the number of unique neuron positions that appear in at least one of the $n$ top-$k$ lists as $z$, the value of $z$ will range between $k$ and $m$.
$S_{l}$ $=(p_{0},p_{1},\ldots,p_{z-1}),$ (9)
$p\in P_{unique}\text{ and }C(p_{0})\geq C(p_{1})\geq\ldots\geq C(p_{z-1})$